OpenStudy (anonymous):

(cos x+1)(2cos^2 x-3cos x-2)? Solve on the interval [0,2pi) A. x=2pi, x=pi/2, x=5pi/4 B. x=pi/6,x=7pi/6 C. x=2pi,x=pi/2,x=pi/3 D. x=pi, x=2pi/3, x=4pi/3

OpenStudy (anonymous):

Is there meant to be a "=0" at the end of that or is it just an expression?

OpenStudy (anonymous):

lol yeah sorry about that there is an =0 at the end.

OpenStudy (anonymous):

Okay, well you can separate into two parts if that's the case: \[(\cos x+1)(2\cos^2 x-3\cos x-2)=0\] \[\cos x+1=0\] or \[2\cos^2 x -3\cos x -2=0\] Are you capable of factorising the one above ^?

OpenStudy (anonymous):

no :( im sorry i dont know how to .

OpenStudy (anonymous):

You just imagine cosx to be x. and you can factorise it like a quadratic.

OpenStudy (anonymous):

I'm sure you must of factorised a quadratic equation before, am I right?

OpenStudy (anonymous):

not sure :/ can you set it up for me please

OpenStudy (anonymous):

Can you factorise this: \[2x^2-3x-2=0\]

OpenStudy (anonymous):

b n

OpenStudy (anonymous):